OpenStudy (nbarrera):

REALLY NEED HELP! WILL MEDAL U! Use the quadratic formula to solve 2y^2 – 3y = 2 A {½, –2} B {–½, –2} C {½, 2} D {–½, 2}

OpenStudy (loser66):

What is the quadratic formula ?

OpenStudy (nbarrera):

isn't up in the description box

OpenStudy (wolf1728):

I know the formula

OpenStudy (loser66):

hahaha.. hence, medal for you @wolf1728

OpenStudy (wolf1728):

a=2 b=-3 c=-2 [x=-b +-sq root(b^2 -4*a*c)] / 2*a

OpenStudy (wolf1728):

[x = --3 +-sq root(9 -4*2*-2)] / 2*2 x = [3 +- sq root (9 +16)] / 4 x = 3 + 5 / 4 x=2 x = 3-5 /4 x= -.5

OpenStudy (wolf1728):

That's D

OpenStudy (wolf1728):

nbarrera - do you agree?

OpenStudy (nbarrera):

yes

OpenStudy (wolf1728):

okay then :-)